A small regional carrier accepted 14 reservations for a particular flight with 10 seats. 5 reservations went to regular customers who will arrive for the flight. Each of the remaining passengers will arrive for the flight with a 46% chance, independently of each other.
(Report answers accurate to 4 decimal places.)

a) Find the probability that overbooking occurs.     
b) Find the probability that the flight has empty seats
Answer #1

Number of seats available = 10

Reservation accepted = 14

Regular customer = 5

Seats left to fill = 10-5 = 5

Reservation left, n = 14-5 =9

Proportion of remaining customer arrival, p = 0.46

Here,

a) Probability of overbooking, P(X>5) = P(6) +P(7) +P(8) +P(9)

b) Probability that flight has empty seats, P(X<5) = 1- P(X>=5)

= 1- P(X=5) - P(X>5)

---------------------------------

If you have any doubt ask in comments.